Shift Summary: 12/13 Owl

Jason LaDue (jladue@cuhep.Colorado.EDU)
Mon, 13 Dec 1999 06:10:26 -0700 (MST)


Shift Summary for owl shift, 12/13/99

On Shift: U. Erkarslan, J. LaDue

Hours of beam available at all intensities: 4.25
Hours of beam available at nominal intensity: 0.00
Hours of E799 data (special runs not included): 3.00
Hours of special runs (laser scans, muon runs...): 0.50

Accounting of lost time:
Accelerator problems with MULAM and Booster Water leak, then
problems getting the beam stop closed (Rad Monitor Trip). Starting and
stopping Runs, KFRENDS, etc.

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:

========================================================================
Accesses during shift:
Access to remove the Regenerator.

========================================================================
Overview of shift:

Beam intensity is not optimal, but at least there is beam.
Typical intensity fluctuated between 7.0E12 and 8.3E12. Lots of SCALER
alarms going off because of poor beam quality.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15209 00:13 Muon run started

00:22 MULAM trips cause beam to go away

02:15 Beam returns at low intensity

02:24 Muon run stopped

02:35 DCGM and DCT0 constants updated

02:40 BobH and JRL go in to pull regenerator

02:55 Out of access

03:09 Problem with A-Loop Interlock, no beam to us

15210 03:00 Pedestal Run started.

03:19 Beam returns.

15211 03:36 BEAM_799 run started, many KTEVANA errors,
Missing FFFF word Stream 6. Put Crate 15
in KFREND list and all works.

04:17 HCC channel 2803 is very ineffecient. Reload
ETOT and HCC it gets better.

94-5 05:41 Pipeline and KTEVANA errors, Init, KFREND, KFREND
reset laser repeater, KFREND, KFREND, reset
repeater, KFREND, KFREND.

06:17 HCC channel 1767 now is really inefficient, page
Jason H, decide to keep an eye on it for now.

========================================================================
Follow up Notes on Previous Shift
SA3 looks fine.
Ed is happy with CsI for now.
DC constants updated.
Muon and Pedestal run taken.

========================================================================
Items to Follow up on Next Shift
CsI 2081 has a large gain drift(~5%), keep an eye on it.
HCC channels 2803 and 1767 were both highly inefficient
during the run, watch to make sure that they aren't getting worse,
experts have been notified.